One number is 7 more than twice the other number. Their sum is 34, find the larger number. How do I work out this problem to fi d the larger number ?
Answer by stanbon(75887)

One number is 7 more than twice the other number. Their sum is 34, find the larger number. How do I work out this problem to find the larger number ?
-----
x = 2y+7
Note:: That means x is larger than y.
-----------------------
x + y = 34
--------
Substitute for "x" and solve for "y"::
34-y = 2y+7
3y = 27
y = 9
-----
x = 2y+7 = 2*9+7 = 25
--------------------------
Cheers,
Stan H.
----------------------